A deck of 10 cards consists of 7 cards that are black on both sides, and 3 cards that are black on one side
and white on the other. 2 cards are selected randomly and placed flat on the table. Calculate the
probability that 1 black and 1 white side are showing.

Answer by Boreal(15235)

There are 20 sides that can be shown, and three are white.
card 1
black 17/20==============white 3/20
card 2=black 16/19; white 3/19=====card 2, black 17/19, white 2/19
The probability will be 17/20*3/19+3/20*17/19=2*51/380=102/380=0.2684